OpenStudy (anonymous):

Solve the problem. Be sure to show all of your work. The wavelength, w, in meters per cycle, of a musical note is given by w = r/f where r is the speed of sound in meters per second and f is the frequency in cycles per second. The speed of sound in air is 344 m/sec. What is the wavelength of a note whose frequency in air is 33 cycles per second? Round to the nearest tenth of a meter per cycle.

sam (.sam.):

w = r/f Look at which of these 3 variable is given: r = 344 m/sec f = 33 cps Substitute these into the original formula w = 344/33 w = 10.4 meters